How do i figure out 2(x+4)-5(x-3)=18
Answer by stanbon(48545)

2(x+4)-5(x-3)=18
2x+8-5x+15=18
-3x+23=18
-3x=-5
x = 5/3
=============
Cheers,
Stan H.